# Graph sensitivity under join and decomposition

### v1

@@ Abstract.
The sensitivity, $\sigma(G)$, of a finite undirected simple graph $G$ is the smallest maximum degree of an induced subgraph on more than the maximum number of independent vertices. Call an indexed family of graphs $G_{n}$ with maximum degree $\Delta(G_{n})\to\infty$ as $n\to\infty$ sensitive if $\sigma(G_{n})\to\infty$, and insensitive otherwise. We describe sensitivity under the join operation and decomposition into stable blocks and construct sensitive and insensitive, primarily non-regular, graph families. We determine the sensitivity explicitly for numerous singly- and doubly-indexed graph families, including certain generalized joins - e.g., complete multipartite graphs and some generalized windmill graphs; general rooted products; and families of corona graphs.
@@ 1. Introduction
The idea of graph sensitivity traces back to work related to Boolean functions and the hypercube graph ([13, 6]), but has been formalized and studied for other graphs following Huang’s clever and concise resolution in [17] of the sensitivity conjecture for Boolean functions [25]. The hypercube graph, $Q_{n}$, which has vertices the $n$-tuples in $\mathbb{Z}_{2}^{n}$ with two vertices adjacent if they differ in exactly one coordinate, is bipartite so its independence number is $\alpha(Q_{n})=\frac{1}{2}2^{n}=2^{n-1}$. Huang proved the smallest maximum degree of an induced subgraph of $Q_{n}$ on $\alpha(Q_{n})+1$ vertices is at least $\sqrt{n}$ and the construction in [6] implies it is equal to $\left\lceil\sqrt{n}\,\right\rceil$. Huang further suggests studying, for other highly symmetric graphs, the smallest maximum degree of an induced subgraph of a graph $G$ on $\alpha(G)+1$ vertices, defined as the sensitivity $\sigma(G)$ in [11] by Garcìa-Marco and Knauer.
@@
Since $Q_{n}$ is the Cayley graph on $\mathbb{Z}_{2}^{n}$ with connection set $S=\{e_{1},\dots,e_{n}\}$ it is natural to investigate the sensitivity of other Cayley graphs. For example, in [1] Alon and Zheng proved $\sigma(\mathrm{Cay}(\mathbb{Z}_{2}^{n},S))\geq\sqrt{|S|}$ for any connection set $S$. In [26], Potechin and Tsang proved a corresponding result for all Cayley graphs of abelian groups using $|V_{G}|/2$ rather than $\alpha(G)$ as the cutoff for the number of vertices, so they prove that $\sqrt{|S|/2}$ provides a lower bound on sensitivity for bipartite graphs. They conjectured that for every Cayley graph $\mathrm{Cay}(G,S)$, the maximum degree of an induced subgraph on more than half the vertices is at least $\sqrt{|S|/2}$, but Lehner and Verret provided an infinite family of counterexamples [23]. Sensitivity and related properties have also been investigated by several authors for the natural generalization of hypercubes to Hamming graphs $H_{n,q}=\mathrm{Cay}(\mathbb{Z}_{q}^{n},\{\pm e_{1},\dots,\pm e_{n}\})$ for $q\geq 3$ (see Theorem 8.1 in [11] and [8, 28, 27, 2]).
@@
The definition of sensitivity of an indexed graph family and three additional infinite families of counterexamples to the Potechin and Tsang conjecture appear in [11] by Garcìa-Marco and Knauer. All of these families have unbounded degree but sensitivity equal to one, which they refer to as insensitive. The Hamming graph families $\{H_{n,q}\}_{n=1}^{\infty}$ for $q\geq 3$ and $\{H_{n,q}\}_{q=3}^{\infty}$ are also insensitive since $\Delta(H_{n,q})=(q-1)n$ and for $q\geq 3$ [28] constructs an induced subgraph of maximum degree one on $\alpha(H_{n,q})+1$ vertices.
@@
In this paper we formalize definitions from [11] and [17], describe graph sensitivity under the join operation and a partition into stable blocks, and apply these results to give general constructions of sensitive and insensitive graph families. We investigate the sensitivity of several explicit, mostly non-regular, families, some of which serve as models for citation, transportation, or neural networks, and find sensitive and insensitive subfamilies occurring within singly- and doubly-indexed families of graphs in a variety of ways. We are aware of two other references that consider sensitivity of nonregular families using a graph operation. Both use a spectral approach as in [17] to analyze Cartesian products: of connected signed graphs one of which is bipartite in [16], and of paths in [30]. More details of the contents are as follows.
@@
In Section 2 after preliminary definitions we present a few examples of sensitive and insensitive graph families, and the definition of $k$-sensitivity, $\sigma_{k}(G)$, of a graph $G$ for $k\geq 1$ as the minimum of the maximum degrees of induced subgraphs on $\alpha(G)+k$ vertices (see the second concluding remark in [17] and Question (5.1) in [30]).
@@ 2. Preliminaries
and we define $\sigma(\overline{K}_{n})=\infty$. An indexed family of graphs $G_{n}$ such that $\Delta(G_{n})\to\infty$ as $n\to\infty$ is sensitive if also $\sigma(G_{n})\to\infty$ and is insensitive if not.
The definition of the sensitivity of a graph and what it means for an indexed graph family to be sensitive or insensitive are motivated by Huang’s result [17] on hypercubes and the discussion in [11]. To compute $\sigma(G)$ it suffices to consider vertex sets with $|S|=\alpha(G)+1$ since the maximum degree can only increase when using a larger vertex set, but all possible such sets must be considered, not just those formed by adding one vertex to a maximum order empty induced subgraph. We define $\sigma(\overline{K}_{n})=\infty$ for convenience in stating Theorem 3.1, but do not consider the family of empty graphs to be sensitive or insensitive. The condition that $\Delta(G_{n})\to\infty$ as $n\to\infty$ distinguishes insensitive graph families from those where $\sigma(G_{n})$ is bounded because $\Delta(G_{n})$ is.
@@
The initial motivating example of a sensitive family is the hypercube family $Q_{n}$, for which $\sigma(Q_{n})\geq\left\lceil\sqrt{n}\,\right\rceil$ [17, 6]. To introduce two simple examples of sensitive non-regular families, recall that the join $G\vee H$ of two graphs $G$ and $H$ is the disjoint union of the two graphs together with all possible edges between them. A natural and useful first example is star graphs.
@@
For any nonempty graph $G$ and for any integer $k\leq|V_{G}|-\alpha(G)$ , define the $k$-sensitivity of $G$ as $\sigma_{k}(G)=0$ if $k\leq 0$ , and if $1\leq k\leq|V_{G}|-\alpha(G)$ as
@@
For an empty graph define
$\sigma_{k}(\overline{K}_{n})\coloneqq\begin{cases}0&\text{if $k\leq 0$ }\\ \infty&\text{if $1\leq k\leq n$.}\end{cases}$
Note that $\sigma_{1}(G)=\sigma(G)$ for all nonempty $G$ and that $\sigma_{k}(G)=\Delta_{\beta}(G)$ for $\beta=\frac{\alpha+k}{|V_{G}|}$ when written in the notation of [11]. We define $\sigma_{k}(\overline{K}_{n})$ to be infinite for $k>1$ as well in order to state Theorem 3.1 uniformly for all graphs.
@@
For Hamming graphs with $q\geq 3$ the sensitivity $\sigma(H_{n,q})=1$ (see [8, 28, 27]). In [27] Potechin and Tsang find additional values of $k>1$ such that $\sigma_{k}(H_{n,3})=1$.
@@ 3.1. Join of two graphs
Let $G_{1}$ and $G_{2}$ be graphs with independence numbers $\alpha_{1}\geq\alpha_{2}$. Then $\alpha(G_{1}\vee G_{2})=\alpha_{1}$. The idea of the proof is to determine the quantities
$a_{j}\coloneqq\min\{\Delta(H)\colon\text{$H$ an induced subgraph with $|V_{H}|=\alpha_{1}+1$ and $|V_{H\cap G_{2}}|=j$}\}$
for $0\leq j\leq\min\{|V_{2}|,\alpha_{1}+1\}$ , whenever they are defined. We may then conclude that $\sigma(G_{1}\vee G_{2})$ is the minimum of these $a_{j}$.
@@
$a_{0}=\min\{\Delta(H)\colon|V_{H}|=\alpha_{1}+1,\ |V_{H\cap G_{2}}|=0\}$
is by definition $\sigma(G_{1})$. Observe that the minimum $a_{0}$ is undefined if $G_{1}$ is empty but by Definition 2.7 we use $a_{0}=\sigma(G_{1})$ regardless.
Next, consider $1\leq j\leq\alpha_{2}$. Since all $j$ vertices in $G_{2}$ are joined to the remaining $\alpha_{1}+1-j$ vertices in $G_{1}$, to obtain a subgraph $H$ with $j$ vertices in $G_{2}$ that achieves the minimum $a_{j}$ the vertices in $G_{1}$ and $G_{2}$ should be taken to be independent and
@@
Finally, suppose that $\alpha_{2}+1\leq j\leq\min\{|V_{2}|,\alpha_{1}\}$. Note that if $G_{2}$ is empty, there is no such $j$. So assume that $G_{2}$ is not empty. In this case, an induced subgraph $H$ of $G_{1}\vee G_{2}$ on $\alpha_{1}+1$ vertices with $|V_{H}\cap V_{2}|=j$ must have at least one edge in $G_{2}$. Thus, in order to find $a_{j}$ we must consider the sensitivity $\sigma_{j-\alpha_{2}}(G_{2})$.
If $|V_{2}|\leq\alpha_{1}+1$, then for $\alpha_{2}+1\leq j\leq|V_{2}|$, let $H$ be an induced subgraph with $j$ vertices in $G_{2}$ that satisfies the following conditions:
  1. (i)
    $H\cap G_{1}$ is empty, and
  2. (ii)
    $H\cap G_{2}$ achieves the sensitivity $\sigma_{j-\alpha_{2}}(G_{2})$.
For every induced subgraph $H^{\prime}$ with $j$ vertices in $G_{2}$ that fails to satisfy either of these properties $\Delta(H^{\prime})\geq\Delta(H)$. Thus
@@
If $\alpha_{1}+1\leq|V_{2}|$, then for $\alpha_{2}+1\leq j<\alpha_{1}+1$, the same argument shows that
@@
For $j=\alpha_{1}+1\leq|V_{2}|$, every induced subgraph on $\alpha_{1}+1$ vertices has all vertices in $G_{2}$. In this case, $a_{j}=\sigma_{j-\alpha_{2}}(G_{2})$.
Taking the minimum of all the quantities $a_{j}$, and noting that Definition 2.7 means that when $G_{2}$ is empty the $a_{j}$ with $\alpha_{2}+1\leq j\leq\min\{|V_{2}|,\alpha_{1}\}$ are disregarded, results in the given values for $\sigma(G_{1}\vee G_{2})$. ∎
When $G_{1}$ and $G_{2}$ are both empty, the expressions in Theorem 3.1 yield an explicit formula for the sensitivity of the complete bipartite graphs $K_{m,n}=\overline{K}_{m}\vee\overline{K}_{n}$ with $m\leq n$. This amounts to splitting $n+1$ vertices as evenly as possible between the two empty graphs, except that when $m\leq\left\lfloor\frac{n+1}{2}\right\rfloor$ it is optimal to include all $m$ vertices in $\overline{K}_{m}$.
@@
If $G=\overline{K}_{m}$, then both  Theorem 3.1 (1) and Theorem 3.1 (2) yield
@@ 3.2. Joining an empty graph to a nonempty graph
We refer to the join of $G$ and the empty graph on $n$ vertices as the $n$-cone of $G$. In this section, we use Example 2.8 to restate Theorem 3.1 in the case where only one of the graphs is empty. This yields more explicit expressions for the sensitivity of the $n$-cone of $G$, allowing exploration of the sensitivity for an array of graph families, including generalized joins like complete multipartite graphs and certain generalized windmill graphs.
@@
This follows directly from Theorem 3.1 by comparing $n=\alpha(\overline{K}_{n})$ with $\alpha(G)$ to determine whether $G$ or $\overline{K}_{n}$ acts as $G_{1}$ in that statement.
In cases (1) and (2), $\alpha(\overline{K}_{n})>\alpha(G)$, so $\overline{K}_{n}$ acts as $G_{1}$ and $G$ as $G_{2}$. The statements in these cases follow from the corresponding parts of Theorem 3.1, noting that when $j\leq\alpha_{2}$ since $\sigma_{j-\alpha_{2}}(G)=0$ then $c_{j}=b_{j}$.
@@
We consider both $\overline{K}_{m}\vee G_{n}$ for $G_{n}=P_{n}$, referred to as agave graphs in [9], and $G_{n}=C_{n}$, referred to as generalized wheel graphs in [4]. We take $m\geq 3$ and $n\geq 4$ or $n\geq 5$ respectively in order to avoid anomalies in the maximum degree and sensitivity for small graphs. With these assumptions on $m$ and $n$ we have
@@
$\sigma(K_{n_{1}}^{\prime}\vee\cdots\vee K_{n_{k}}^{\prime})=\sigma(K_{n_{1}}^{\prime}\vee K_{n_{2}}^{\prime})=\begin{cases}n_{1}+1-n_{2}&\text{ if $n_{2}\leq\left\lfloor\frac{n_{1}+1}{2}\right\rfloor$\,,}\\ \left\lceil\frac{n_{1}+1}{2}\right\rceil&\text{ if $n_{2}>\left\lfloor\frac{n_{1}+1}{2}\right\rfloor$.}\end{cases}$
@@
The generalized join is also referred to as the joined union [5] or Zykov sum [24]. Many common constructions can be realized as other special cases of generalized joins. For example, $K_{2}[G_{1},G_{2}]=G_{1}\vee G_{2}$ is the standard join, $G[H,\dots,H]=G\circ H$ is the lexicographic product or graph composition of $G$ and $H$, and the complete multipartite graphs in Proposition 3.11 are generalized joins of empty graphs over a complete graph, i.e., of the form $K_{m}[\overline{K}_{n_{1}},\dots,\overline{K}_{n_{m}}]$.
The windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$ in Example 3.6, which can be viewed as the generalized join $S_{m+1}[K_{1},K_{n},\dots,K_{n}]$, were used in [9] to explore divergence between often-used transitivity indices and proposed as appropriate models for citation and collaboration networks that exhibit similar divergence properties. In order to extend and expand these results, in [18] Kooij considered both what he called type I generalized windmill graphs $K_{m}\vee nH=S_{n+1}[K_{m},H,\dots,H]$ and type II generalized windmill graphs $\overline{K}_{m}\vee nH=S_{n+1}[\overline{K}_{m},H,\dots,H]$, as well as a type III graph that is actually a corona - see Definition 4.17. The $W_{m,n}$ and Types I and II in [18] are all subsumed as particular instances of $G[H_{0},H_{1},\dots,H_{n}]$ with $G=S_{n+1}$. The generalized join or joined union $G[H_{1},\dots,H_{n}]$ for any graph $G$ of order $n$ is described as a fourth family of generalized windmill graphs in [24], where it is noted they more closely relate to real-life transportation networks, and in [5] it is pointed out that the graphs $H_{1},\dots,H_{n}$ can be considered to describe the layers of a multilayer network with $G$ controlling how the layers are joined. We consider a few cases of generalized joins for which we can determine sensitivity.
@@
For arbitrary $H_{1},\dots,H_{n}$ Theorem 3.1 still applies successively, but even if the $H_{i}$ were ordered to be monotonic in $\alpha_{i}$, computing $\sigma(K_{m}[H_{1},\dots,H_{m}])$ could require $k$-sensitivities of some $H_{i}$ and switching between using Theorem 3.1 (1) and (2) if there is not a consistent comparison between the larger of the independence numbers and the order of the graph with smaller independence number. Being able to fully compute $\sigma(K_{m}[H_{1},\dots,H_{m}])$ would address the series decomposition, which describes the partition of $G^{\prime}$ into connected components $H^{\prime}_{1},\dots,H^{\prime}_{m}$ (see Section 5).
@@
For the generalized windmill graphs of type I in [18], by Corollary 3.4 and Example 3.14 we have

### v2

@@ Abstract.
The sensitivity, $\sigma(G)$, of a finite undirected simple graph $G$ is the smallest maximum degree of an induced subgraph on more than the maximum number of independent vertices. Call an indexed family of graphs $G_{n}$ with maximum degree $\Delta(G_{n})\to\infty$ as $n\to\infty$ sensitive if $\sigma(G_{n})\to\infty$, and insensitive otherwise. These definitions have their roots in Huang’s resolution of the Sensitivity Conjecture for Boolean functions via determining sensitivity for the hypercube graphs and subsequent investigations of other Cayley graph families. Here we describe sensitivity under the join operation and decomposition into stable blocks and construct sensitive and insensitive, primarily non-regular, graph families. We determine the sensitivity explicitly for numerous singly- and doubly-indexed graph families, including certain generalized joins - e.g., complete multipartite graphs and some generalized windmill graphs; general rooted products; and families of corona graphs.
@@ 1. Introduction
The idea of graph sensitivity traces back to work related to Boolean functions and the hypercube graph ([Gotsman-Linial-92, CFGS-88]), but has been formalized and studied for other graphs following Huang’s clever and concise resolution in [Huang-19] of the sensitivity conjecture for Boolean functions [Nisan-Szegedy-94]. The hypercube graph, $Q_{n}$, which has vertices the $n$-tuples in $\mathbb{Z}_{2}^{n}$ with two vertices adjacent if they differ in exactly one coordinate, is bipartite so its independence number is $\alpha(Q_{n})=\frac{1}{2}2^{n}=2^{n-1}$. Huang proved the smallest maximum degree of an induced subgraph of $Q_{n}$ on $\alpha(Q_{n})+1$ vertices is at least $\sqrt{n}$ and the construction in [CFGS-88] implies it is equal to $\left\lceil\sqrt{n}\,\right\rceil$. Huang further suggests studying, for other highly symmetric graphs, the smallest maximum degree of an induced subgraph of a graph $G$ on $\alpha(G)+1$ vertices, defined as the sensitivity $\sigma(G)$ in [GarciaMarco-Knauer-22] by Garcìa-Marco and Knauer.
@@
Since $Q_{n}$ is the Cayley graph on $\mathbb{Z}_{2}^{n}$ with connection set $S=\{e_{1},\dots,e_{n}\}$ it is natural to investigate the sensitivity of other Cayley graphs. For example, in [Alon-Zheng-20] Alon and Zheng proved $\sigma(\mathrm{Cay}(\mathbb{Z}_{2}^{n},S))\geq\sqrt{|S|}$ for any connection set $S$. In [Potechin-Tsang-20], Potechin and Tsang proved a corresponding result for all Cayley graphs of abelian groups using $|V_{G}|/2$ rather than $\alpha(G)$ as the cutoff for the number of vertices, so they prove that $\sqrt{|S|/2}$ provides a lower bound on sensitivity for bipartite graphs. They conjectured that for every Cayley graph $\mathrm{Cay}(G,S)$, the maximum degree of an induced subgraph on more than half the vertices is at least $\sqrt{|S|/2}$, but Lehner and Verret provided an infinite family of counterexamples [Lehner-Verret-20]. Sensitivity and related properties have also been investigated by several authors for the natural generalization of hypercubes to Hamming graphs $H_{n,q}=\mathrm{Cay}(\mathbb{Z}_{q}^{n},\{\pm e_{1},\dots,\pm e_{n}\})$ for $q\geq 3$ (see Theorem 8.1 in [GarciaMarco-Knauer-22] and [Dong-21, Tandya-22, Potechin-Tsang-24, AFGMK-25]).
@@
The definition of sensitivity of an indexed graph family and three additional infinite families of counterexamples to the Potechin and Tsang conjecture appear in [GarciaMarco-Knauer-22] by Garcìa-Marco and Knauer. All of these families have unbounded degree but sensitivity equal to one, which they refer to as insensitive families. The Hamming graph families $\{H_{n,q}\}_{n=1}^{\infty}$ for $q\geq 3$ and $\{H_{n,q}\}_{q=3}^{\infty}$ are also insensitive since $\Delta(H_{n,q})=(q-1)n$ and for $q\geq 3$ [Tandya-22] constructs an induced subgraph of maximum degree one on $\alpha(H_{n,q})+1$ vertices.
@@
In this paper we formalize definitions from [GarciaMarco-Knauer-22] and [Huang-19], describe graph sensitivity under the join operation and a partition into stable blocks, and apply these results to give general constructions of sensitive and insensitive graph families. We investigate the sensitivity of several explicit, mostly non-regular, families, some of which serve as models for citation, transportation, or neural networks, and find sensitive and insensitive subfamilies occurring within singly- and doubly-indexed families of graphs in a variety of ways. We are aware of two other references that consider sensitivity of nonregular families using a graph operation. Both use a spectral approach as in [Huang-19] to analyze Cartesian products: of connected signed graphs one of which is bipartite in [HLL-21], and of paths in [ZH-24]. More details of the contents are as follows.
@@
In Section 2 after preliminary definitions we present a few examples of sensitive and insensitive graph families, and the definition of $k$-sensitivity, $\sigma_{k}(G)$, of a nonempty graph $G$ for $1\leq k\leq|V_{G}|-\alpha(G)$ as the minimum of the maximum degrees of induced subgraphs on $\alpha(G)+k$ vertices (see the second concluding remark in [Huang-19] and Question (5.1) in [ZH-24]).
@@ 2. Preliminaries
and we define $\sigma(\overline{K}_{n})=n$. [^1] We thank Martin Rubey for proposing $\sigma(\overline{K}_{n})=n$ as more natural than $\infty$ since then adding a dominant vertex to an empty graph preserves sensitivity (see Example 2.4), and for adding graph sensitivity to the database findstat.org. An indexed family of graphs $G_{n}$ such that $\Delta(G_{n})\to\infty$ as $n\to\infty$ is sensitive if also $\sigma(G_{n})\to\infty$ and is insensitive if not.
The definition of the sensitivity of a graph and what it means for an indexed graph family to be sensitive or insensitive are motivated by Huang’s result [Huang-19] on hypercubes and the discussion in [GarciaMarco-Knauer-22]. To compute $\sigma(G)$ for a nonempty graph it suffices to consider vertex sets with $|S|=\alpha(G)+1$ since the maximum degree can only increase when using a larger vertex set, but all possible such sets must be considered, not just those formed by adding one vertex to a maximum order empty induced subgraph. Defining $\sigma(\overline{K}_{n})$ to be $n$ allows us to state Theorem 3.1 uniformly for all graphs. We do not consider the family of empty graphs to be sensitive or insensitive and in general the condition that $\Delta(G_{n})\to\infty$ as $n\to\infty$ distinguishes insensitive graph families from those where $\sigma(G_{n})$ is bounded because $\Delta(G_{n})$ is bounded.
@@
The initial motivating example of a sensitive family is the hypercube family $Q_{n}$, for which $\sigma(Q_{n})\geq\left\lceil\sqrt{n}\,\right\rceil$ [Huang-19, CFGS-88]. To introduce two simple examples of sensitive non-regular families, recall that the join $G\vee H$ of two graphs $G$ and $H$ is the disjoint union of the two graphs together with all possible edges between them. A natural and useful first example is star graphs.
@@
For any nonempty graph $G$ and any integer $1\leq k\leq|V_{G}|-\alpha(G)$ , define the $k$-sensitivity of $G$ as
@@
For an empty graph define $\sigma_{k}(\overline{K}_{n})=n$ for all $k\geq 1$ and for any graph $G$ define $\sigma_{k}(G)=0$ if $k\leq 0$.
Note that for all graphs $\sigma_{1}(G)=\sigma(G)$ and for all nonempty graphs we have $\sigma_{k}(G)=\Delta_{\beta}(G)$ for $\beta=\frac{\alpha(G)+k}{|V_{G}|}$ when written in the notation of [GarciaMarco-Knauer-22].
@@
For Hamming graphs with $q\geq 3$ the sensitivity $\sigma(H_{n,q})=1$ (see [Dong-21, Tandya-22, Potechin-Tsang-24]). In [Potechin-Tsang-24] Potechin and Tsang find additional values of $k>1$ such that $\sigma_{k}(H_{n,3})=1$.
@@ 3.1. Join of two graphs
Let $G_{1}$ and $G_{2}$ be graphs with vertex sets $V_{1}$ and $V_{2}$ and independence numbers $\alpha_{1}\geq\alpha_{2}$. Hence $\alpha(G_{1}\vee G_{2})=\alpha_{1}$. We will determine the quantities
$a_{j}\coloneqq\min\{\Delta(H)\colon\text{$H$ an induced subgraph with $|V_{H}|=\alpha_{1}+1$ and $|V_{H}\cap V_{2}|=j$}\}$
with $0\leq j\leq\min\{|V_{2}|,\alpha_{1}+1\}$, whenever they are defined. Then $\sigma(G_{1}\vee G_{2})$ will be the minimum of these $a_{j}$.
@@
$a_{0}=\min\{\Delta(H)\colon|V_{H}|=\alpha_{1}+1,\ |V_{H}\cap V_{2}|=0\}$
is by definition $\sigma(G_{1})$. Observe that if $G_{1}$ is empty, then $a_{0}$ is undefined, but in this case by Definition 2.7 $a_{1}=|V_{1}|=\sigma(G_{1})$, so we use $a_{0}=\sigma(G_{1})$ in this case as well.
Next, consider $1\leq j\leq\alpha_{2}$. Since all $j$ vertices of $H$ in $G_{2}$ are joined to the remaining $\alpha_{1}+1-j$ vertices of $H$ in $G_{1}$, to minimize $\Delta(H)$ we use $j$ independent vertices in $G_{2}$ and $\alpha_{1}+1-j$ independent vertices in $G_{1}$. Thus for $1\leq j\leq\alpha_{2}$ and such a subgraph $H$,
@@
Finally, suppose that $\alpha_{2}+1\leq j\leq\min\{|V_{2}|,\alpha_{1}+1\}$. Note that if $G_{2}$ is empty then $|V_{2}|=\alpha_{2}\leq\alpha_{1}$ and there is no such $j$. So assume that $G_{2}$ is not empty. In this case, an induced subgraph $H$ of $G_{1}\vee G_{2}$ on $\alpha_{1}+1$ vertices with $|V_{H}\cap V_{2}|=j$ must have at least one edge in $G_{2}$. Thus, in order to find $a_{j}$ we must consider the sensitivity $\sigma_{j-\alpha_{2}}(G_{2})$.
If $|V_{2}|<\alpha_{1}+1$ then there will be at least one vertex in $G_{1}$ and to minimize $\Delta(H)$ the vertices of $H$ in $G_{2}$ must be chosen so that $\Delta(H\cap G_{2})=\sigma_{j-\alpha_{2}}(G_{2})$, i.e., to have the subgraph induced by $H$ within $G_{2}$ be of smallest maximum degree. Then for $\alpha_{2}+1\leq j\leq|V_{2}|$ and such a subgraph $H$,
@@
If $|V_{2}|\geq\alpha_{1}+1$, then for $\alpha_{2}+1\leq j\leq\alpha_{1}$ the same argument shows that
@@
But when $j=\alpha_{1}+1\leq|V_{2}|$, then every induced subgraph used in computing $a_{j}=a_{\alpha_{1}+1}$ has all vertices in $G_{2}$ and in this case, $a_{\alpha_{1}+1}=\sigma_{\alpha_{1}+1-\alpha_{2}}(G_{2})$.
Taking the minimum of all the quantities $a_{j}$ and introducing the notation $b_{j}$ and $c_{j}$ results in the given values for $\sigma(G_{1}\vee G_{2})$. ∎
It follows from the proof that in Theorem 3.1 (1) and (2), when $G_{1}$ is empty $b_{1}=|G_{1}|=\sigma(G_{1})$ by Definition 2.7 and $\sigma(G_{1})$ may either be included or omitted, and when $G_{2}$ is empty the $c_{j}$ may be omitted. This combined with Lemma 2.12 means that when $G_{1}$ and $G_{2}$ are both empty the expressions in Theorem 3.1 yield an explicit formula for the sensitivity of the complete bipartite graphs $K_{m,n}=\overline{K}_{m}\vee\overline{K}_{n}$ with $m\leq n$. This can also be found directly and amounts to splitting $n+1$ vertices as evenly as possible between the two empty graphs, except that when $m\leq\left\lfloor\frac{n+1}{2}\right\rfloor$ it is optimal to include all $m$ vertices in $\overline{K}_{m}$.
@@
If $G=\overline{K}_{m}$, then both Theorem 3.1 (1) and (2) yield
@@ 3.2. Joining an empty graph to a nonempty graph
We refer to the join of $G$ and the empty graph on $n$ vertices as the $n$-cone of $G$. In this section, we use Definition 2.7 to restate Theorem 3.1 in the case where only one of the graphs is empty. This yields more explicit expressions for the sensitivity of the $n$-cone of $G$, allowing exploration of the sensitivity for an array of graph families, including generalized joins like complete multipartite graphs and certain generalized windmill graphs.
@@
This follows directly from Theorem 3.1 by comparing $n=\alpha(\overline{K}_{n})$ with $\alpha(G)$ to determine whether $G$ or $\overline{K}_{n}$ serves as $G_{1}$ in that statement.
In cases (1) and (2), $\alpha(\overline{K}_{n})>\alpha(G)$, so use $\overline{K}_{n}$ as $G_{1}$ and $G$ as $G_{2}$. The statements in these cases follow from the corresponding parts of Theorem 3.1, noting that when $j\leq\alpha_{2}$ since $\sigma_{j-\alpha_{2}}(G)=0$ then $b_{j}=c_{j}$, and that $c_{1}=\max\{1,n+\sigma_{1-\alpha(G)}(G)\}=\max\{1,n\}=n$, so $\sigma(\overline{K}_{n})=n$ may be omitted when taking the minimum.
@@
We consider both $\overline{K}_{m}\vee G_{n}$ for $G_{n}=P_{n}$, referred to as agave graphs in [Estrada-16], and $G_{n}=C_{n}$, referred to as generalized wheel graphs in [Buckley-Harary-88]. We take $m\geq 3$ and $n\geq 4$ or $n\geq 5$ respectively in order to avoid anomalies in the maximum degree and sensitivity for small graphs. With these assumptions on $m$ and $n$ we have
@@
$\sigma(\overline{K}_{n_{1}}\vee\cdots\vee\overline{K}_{n_{k}})=\sigma(\overline{K}_{n_{1}}\vee\overline{K}_{n_{2}})=\begin{cases}n_{1}+1-n_{2}&\text{ if $n_{2}\leq\left\lfloor\frac{n_{1}+1}{2}\right\rfloor$\,,}\\ \left\lceil\frac{n_{1}+1}{2}\right\rceil&\text{ if $n_{2}>\left\lfloor\frac{n_{1}+1}{2}\right\rfloor$.}\end{cases}$
@@
The generalized join is also referred to as the joined union [Chudnovsky-24] or Zykov sum [Liao-Aziz-Alaoui-Hou-22]. Many common constructions can be realized as other special cases of generalized joins. For example, $K_{2}[G_{1},G_{2}]=G_{1}\vee G_{2}$ is the standard join, $G[H,\dots,H]=G\circ H$ is the lexicographic product or graph composition of $G$ and $H$, and the complete multipartite graphs in Proposition 3.11 are generalized joins of empty graphs over a complete graph, i.e., of the form $K_{m}[\overline{K}_{n_{1}},\dots,\overline{K}_{n_{m}}]$.
The windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$ in Example 3.6, which can be viewed as the generalized join $S_{m+1}[K_{1},K_{n},\dots,K_{n}]$, were used in [Estrada-16] to explore divergence between often-used transitivity indices and proposed as appropriate models for citation and collaboration networks that exhibit similar divergence properties. In order to extend and expand these results, in [Kooij-19] Kooij considered both what he called type I generalized windmill graphs $K_{m}\vee nH=S_{n+1}[K_{m},H,\dots,H]$ and type II generalized windmill graphs $\overline{K}_{m}\vee nH=S_{n+1}[\overline{K}_{m},H,\dots,H]$, as well as a type III graph that is actually a corona - see Definition 4.17. The $W_{m,n}$ and Types I and II in [Kooij-19] are all subsumed as particular instances of $G[H_{0},H_{1},\dots,H_{n}]$ with $G=S_{n+1}$. The generalized join or joined union $G[H_{1},\dots,H_{n}]$ for any graph $G$ of order $n$ is described as a fourth family of generalized windmill graphs in [Liao-Aziz-Alaoui-Hou-22], where it is noted they more closely relate to real-life transportation networks, and in [Chudnovsky-24] it is pointed out that the graphs $H_{1},\dots,H_{n}$ can be considered to describe the layers of a multilayer network with $G$ controlling how the layers are joined. We consider a few cases of generalized joins for which we can determine sensitivity.
@@
For arbitrary $H_{1},\dots,H_{n}$ Theorem 3.1 still applies successively, but even if the $H_{i}$ were ordered to be monotonic in $\alpha_{i}$, computing $\sigma(K_{m}[H_{1},\dots,H_{m}])$ could require $k$-sensitivities of some $H_{i}$ and switching between using Theorem 3.1 (1) and (2) if there is not a consistent comparison between the larger of the independence numbers and the order of the graph with smaller independence number. Being able to fully compute $\sigma(K_{m}[H_{1},\dots,H_{m}])$ would address the series decomposition, which describes the partition of $\overline{G}$ into connected components $\overline{H}_{1},\dots,\overline{H}_{m}$ (see Section 5).
@@
For the generalized windmill graphs of type I in [Kooij-19], by Corollary 3.4 and Example 3.14 we have